OpenStudy (anonymous):

If a+b+c=0 and a^2+b^2+c^2=22, what is a^4+b^4+c^4?

OpenStudy (jack1):

you could assume that a = b+c so if a + b + c = 0 = true then a^2+b^2+c^2=22 so as a = b+c (b+c)^2 +b^2+c^2=22

OpenStudy (anonymous):

then

OpenStudy (jack1):

(b+c)^2 +b^2+c^2=22 b^2 +2bc + c^2 +b^2 + c^2 = 22 2b^2+ 2c^2 + 2bc = 22

OpenStudy (anonymous):

Then?

OpenStudy (jack1):

then reduce, so b = something + c sub back into eqn 1 solve simultaneous equations

OpenStudy (anonymous):

help me for that

OpenStudy (jack1):

factor this: 2b^2+ 2c^2 + 2bc = 22 in terms of b

OpenStudy (jack1):

first: divide both sides by 2

OpenStudy (jack1):

then try quadratic solution after rearranging for = 0...?

OpenStudy (jack1):

2b^2+ 2c^2 + 2bc = 22 so 2b^2+ 2c^2 + 2bc -22 = 0 so b^2+ c^2 + bc -11 = 0 use quadratic solution b = .5 + or - sqrt ((44 - 3c^2) - c)

OpenStudy (jack1):

sub back in, solve for c so c = sqrt (11/3) therefore b = -2 (sqrt (11/3))

OpenStudy (anonymous):

so what is the answer

OpenStudy (jack1):

so a must = c = sqrt (11/3) (as a+b+c = 0, in this case a+c = b as we solved for b first)

OpenStudy (jack1):

now just put each to the power of 4 to find ur answer

OpenStudy (anonymous):

answer?

OpenStudy (jack1):

answer is a^4+b^4+c^4 where: a = sqrt (11/3) b = -2 (sqrt (11/3)) c = sqrt (11/3) remember that sqrt a = a^(1/2)

OpenStudy (anonymous):

what is the value of a^4+b^4+c^4

OpenStudy (jack1):

u work it out, i've given you the values of a and b and c

OpenStudy (jack1):

im here to help you do your work, not do it for you how bout u give it a try, and i'll tell u if you're right?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

1210/9

OpenStudy (jack1):

no.

OpenStudy (jack1):

try approximating the fractions into just numbers, less confusion that way what's 11/3 =

OpenStudy (anonymous):

3.66

OpenStudy (jack1):

then take the sqr root of that

OpenStudy (anonymous):

0.6

OpenStudy (jack1):

no, 1.915

OpenStudy (anonymous):

oh i see then?

OpenStudy (jack1):

a = 1.915 so a^4 = 1.915 x 1.915 x 1.915 x 1.915 =???

OpenStudy (anonymous):

13.44

OpenStudy (jack1):

perfect now c = 1.915 so c^4 =???

OpenStudy (anonymous):

13.44

OpenStudy (jack1):

spot on, now b = -2 x sqrt 11/3 =-2 x 1.915 =?

OpenStudy (anonymous):

-3.830

OpenStudy (jack1):

close enough, i got -3.8297 so b = -3.830 therefore b^4 = -3.830 x -3.830 x -3.830 x -3.830 = ???

OpenStudy (anonymous):

215.17

OpenStudy (jack1):

yep, now add them together

OpenStudy (anonymous):

thank you

OpenStudy (jack1):

welcome what was ur final answer?

OpenStudy (anonymous):

242.( )( )

OpenStudy (jack1):

yep, the . is only there because of rounding and approximating, whole answer is 242 even

OpenStudy (asnaseer):

@Jack1 - how can you assume a = b + c in your first step?

OpenStudy (jack1):

sorry, a = -b-c

OpenStudy (jack1):

if a + b + c = 0 ... my bad

OpenStudy (asnaseer):

ok - that makes more sense

OpenStudy (jack1):

and i ended up choosing b to equal (-c + -a) instead anyway

OpenStudy (jack1):

as my calcs ended up with: a = sqrt (11/3) b = -2 (sqrt (11/3)) c = sqrt (11/3)

OpenStudy (asnaseer):

here is an alternative method to solve this. given:\[a+b+c=0\tag{1}\]\[a^2+b^2+c^2=22\tag{2}\]then:\[\begin{align} (a+b+c)^2&=a^2+b^2+c^2+2(ab+bc+ca)=22+2(ab+bc+ca)\\ \therefore 0&=22+2(ab+bc+ca)\\ \therefore ab+bc+ca&=-11\tag{3}\\ (ab+bc+ca)^2&=a^2b^2+b^2c^2+c^2a^2+2(ab^2c+a^2bc+abc^2)\\ &=a^2b^2+b^2c^2+c^2a^2+2abc(a+b+c)\\ \therefore (-11)^2&=a^2b^2+b^2c^2+c^2a^2+0\\ \therefore a^2b^2+b^2c^2+c^2a^2&=121\tag{4} \end{align}\]then finally we can get:\[\begin{align} (a^2+b^2+c^2)^2&=a^4+b^4+c^4+2(a^2b^2+b^2c^2+c^2a^2)\\ \therefore (22)^2&=a^4+b^4+c^4+2(121)\\ \therefore a^4+b^4+c^4&=22^2-242=484-242=242\\ \end{align}\]

OpenStudy (jack1):

much cleaner